# Physical insights from the aspect ratio dependence of turbulence in negative triangularity plasmas

### v1

@@ VII Conclusions
  • In the second part, we started from a simplified model of the toroidal ITG instability and were able to explain the beneficial effect of NT on confinement for ITG at large and conventional A. Using linear and nonlinear simulations, we showed that the stabilization of ITG in NT comes from faster magnetic drifts with respect to a resonance condition set by the diamagnetic drift, and from stronger FLR effects. However, even though the beneficial effect of NT is still present at small A in ITG regimes, the physical picture developed in this paper does not apply as many other physical effects are important.
  • In the third part, we performed the first GK simulations of the NT and PT equilibria predicted for SMART. Linear analysis showed that MTMs dominate at ion scales. Moreover, in NT they have larger linear growth rate when compared to PT with the same kinetic parameters. A scan in electron $\beta$ showed that MTMs appear at lower $\beta$ and $R/L_{Te}$ in NT and have larger growth rates. However, when ITG is the dominant type of instability, NT has weaker linear growth rates, which is compatible with the findings of the previous sections. Finally, a scan of the magnetic shear showed that large values are crucial to drive MTMs at large $\beta$ and electron temperature gradients. A triple scan in ion temperature gradient, electron temperature gradient and magnetic shear showed that by lowering the magnetic shear and the ratio of electron temperature gradient to ion temperature gradient it is possible to extend the space where SMART NT option can be operated in ITG-dominated regime and exploit the beneficial effect of NT.
@@ Appendix A Linear simulations to establish turbulent regime
In this appendix, we show the results of the linear simulations performed for the different scenarios shown in figure 3 to determine the turbulent regime as the aspect ratio is changed. Figure 25 shows the growth rates and the frequencies as functions of the binormal wavenumber for the $k_{x}=0$ mode. We see that, as expected, the real TEM and pTEM cases are dominated by TEM turbulence while the pITG-2 and real ITG are dominated by ITG turbulence.
Appendix B Field-aligned coordinates and GENE
@@
It is clear that this is not an orthogonal coordinate system, but a curvilinear one. Hence, the standard algebra for curvilinear systems can be applied [32]. We can define the metric coefficients $g^{ij}$ and $g_{ij}$, which define the differential arc length along a curve in the curvilinear system
@@
which is related to the metric coefficients through [32]
@@
Appendix C Derivation of magnetic drift frequency geometric dependence
@@
Appendix D NT ITG physical picture applied to elongation and magnetic shear scans
@@ D.1 Elongation scan
Figure 26: Ratio of NT heat flux over heat flux of the PT counterpart as a function of the plasma elongation. The dashed black lines highlight a ratio of 1, where NT and PT have the same confinement.
@@
D.2 Magnetic shear scan
We performed the same study as in the previous section, but with the magnetic shear $\hat{s}$ instead of elongation. Magnetic shear $\hat{s}$ was changed from large positive to negative values of shear, in the range $\hat{s}=[-0.5,2.0]$. Figure 28 shows the ratio between the nonlinear heat flux from the NT cases and the PT cases as a function of the magnetic shear. We observe a monotonic decrease of the ratio with $\hat{s}$. We could not find a value of magnetic shear for which NT becomes detrimental, but NT becomes less beneficial at lower $\hat{s}$. These observations are consistent with [10]. However, here we can try to apply our NT ITG physical picture to explain this trend.

### v2

@@ VII Conclusions
  • In the second part, we performed the first GK simulations of the NT and PT equilibria predicted for SMART. Linear analysis showed that MTMs dominate at ion scales. Moreover, in NT they have larger linear growth rate when compared to PT with the same kinetic parameters. A scan in electron $\beta$ showed that MTMs appear at lower $\beta$ and $R/L_{Te}$ in NT and have larger growth rates. However, when ITG is the dominant type of instability, NT has weaker linear growth rates, which is compatible with the findings of the previous section. Finally, a scan of the magnetic shear showed that large values are crucial to drive MTMs at large $\beta$ and electron temperature gradients. A triple scan in ion temperature gradient, electron temperature gradient and magnetic shear showed that by lowering the magnetic shear and the ratio of electron temperature gradient to ion temperature gradient it is possible to extend the space where SMART NT option can be operated in ITG-dominated regime and exploit the beneficial effect of NT.
  • In the third part, we started from a simplified model of the toroidal ITG instability and were able to explain the beneficial effect of NT on confinement for ITG at large and conventional A. Using linear and nonlinear simulations, we showed that the stabilization of ITG in NT comes from faster magnetic drifts with respect to a resonance condition set by the diamagnetic drift, and from stronger FLR effects. However, even though the beneficial effect of NT is still present at small $A$ in ITG regimes, the physical picture developed in this paper does not apply as many other physical effects are important.
@@ Appendix A Linear simulations to establish turbulent regime
In this appendix, we show the results of the linear simulations performed for the different scenarios shown in figure 2 to determine the turbulent regime as the aspect ratio is changed. Figure 23 shows the growth rates and the frequencies as functions of the binormal wavenumber for the $k_{x}=0$ mode. We see that, as expected, the real TEM and pTEM cases are dominated by TEM turbulence while the pITG-1, pITG-2 and real ITG are dominated by ITG turbulence.
Appendix B Derivation of the condition for the optimum growth rate of an ITG mode
In this appendix we report the derivation that leads to equation (5). Toroidal ITG modes are driven by the magnetic field line curvature and gradients in the magnetic field strength. They can be unstable where the curvature is “bad”, meaning the magnetic field curvature (or equivalently gradient) has a component parallel to the temperature gradient.
We can estimate the characteristic value of this velocity as
$\textbf{v}_{D}=\frac{v_{th}^{2}}{B\Omega}\textbf{b}\wedge\bm{\nabla}B,$ (14)
$v_{E\wedge B}=\frac{\textbf{E}\wedge\textbf{B}}{B^{2}}.$ (15)
One can derive a dispersion relation for the ITG modes, which corresponds to the qualitative physical picture Beer (1994). In gyrofluid theory, it is given by.
$\omega_{D}=\frac{(\tau+b)\eta}{1+3(\tau+b)^{2}}\omega_{*}.$ (19)
This is the optimal condition for the onset of an unstable toroidal ITG.
Appendix C Field-aligned coordinates and GENE
@@
It is clear that this is not an orthogonal coordinate system, but a curvilinear one. Hence, the standard algebra for curvilinear systems can be applied D’haeseleer et al. (1991). We can define the metric coefficients $g^{ij}$ and $g_{ij}$, which define the differential arc length along a curve in the curvilinear system
@@
which is related to the metric coefficients through D’haeseleer et al. (1991)
@@
Appendix D Derivation of magnetic drift frequency geometric dependence
@@
Appendix E NT ITG physical picture applied to elongation and magnetic shear scans
@@ E.1 Elongation scan
Figure 25: Ratio of NT heat flux over heat flux of the PT counterpart as a function of the plasma elongation. The dashed black line highlight a ratio of 1, where NT and PT have the same confinement.
@@
E.2 Magnetic shear scan
We performed the same study as in the previous section, but with the magnetic shear $\hat{s}$ instead of elongation. Magnetic shear $\hat{s}$ was changed from large positive to negative values of shear, in the range $\hat{s}=[-0.5,2.0]$. Figure 27 shows the ratio between the nonlinear heat flux from the NT cases and the PT cases as a function of the magnetic shear. We observe a monotonic decrease of the ratio with $\hat{s}$. We could not find a value of magnetic shear for which NT becomes detrimental, but NT becomes less beneficial at lower $\hat{s}$. These observations are consistent with Merlo and Jenko (2023). However, here we can try to apply our NT ITG physical picture to explain this trend.